# Multi-Hop Question Generation via Dual-Perspective Keyword Guidance

### v1

@@ Abstract
Maodong Li1, Longyin Zhang2, Fang Kong1thanks: Corresponding authors 1School of Computer Science and Technology, Soochow University, China 2Institute for Infocomm Research, A*STAR, Singapore {20234227022@stu,kongfang@}suda.edu.cnzhang_longyin@i2r.a-star.edu.sg
@@ 3.2 DPKG Framework
where $H^{i}_{doc}$ and $H^{i}_{ans}$ represent document states and answer states, respectively. $H^{i}_{da}$ denotes document-answer states, and [;][;][ ; ] denotes concatenation. Notably, we modified the internal structure of the transformer encoder to ensure that, when encoding the three input components simultaneously, each part remains independent, except for the shared self-attention mechanism. We encode the document and answer separately to facilitate the subsequent computation of answer-aware states. In Figure 2, (b) and (c) are identical, except (c) incorporates an additional vanilla BART encoder Lewis et al. (2020) to encode keywords. Although structurally similar, they are independent because of their distinct tasks. Specifically, (b) generates keywords based on the document and answer, while (c) leverages the keywords to pinpoint essential information snippets related to QA pairs in the document and generate multi-hop questions. Formalizing these processes:
@@ 4.1 Dataset
Table 1: Statistics of datasets, where "Sam." denotes "Sample," "Qes." denotes "Question keyword," and "Doc." denotes "Document keyword."
@@ C.2 Results of Fine-tuned Large Language Models
It can be seen from Table 8 that the performance of Llama in the SF and Full settings is essentially similar. This is expected, as Llama has a large number of parameters and strong contextual understanding ability. For Llama, the text length differences in the SF and Full settings are insufficient to cause significant performance variation. In the SF setting, DPKG still performs well on some metrics. However, in the Full setting, its performance does not match that of Llama — which is expected, given Llama’s significantly larger parameter count. Nevertheless, DPKG still delivers competitive results, further demonstrating the effectiveness of the proposed dual-perspective keywords and the DPKG framework.
@@ Appendix E Generalization to Other Datasets
Table 12: Statistics of MusiQue-2hop datasets, where "Sam." denotes "Sample," "Qes." denotes "Question keyword," and "Doc." denotes "Document keyword."
Model B-4 MTR R-L

### v2

@@ Abstract
Maodong Li1, Longyin Zhang2, Fang Kong1thanks: Corresponding author 1School of Computer Science and Technology, Soochow University, China 2Institute for Infocomm Research, A*STAR, Singapore {20234227022@stu,kongfang@}suda.edu.cnzhang_longyin@i2r.a-star.edu.sg
@@ 3.2 DPKG Framework
where $H^{i}_{doc}$ and $H^{i}_{ans}$ represent document states and answer states, respectively. $H^{i}_{da}$ denotes document-answer states, and [;][;][ ; ] denotes concatenation. Notably, we modify the internal structure of the transformer encoder to ensure that, when encoding the three input components simultaneously, each part remains independent, except for the shared self-attention mechanism. We encode the document and answer separately to facilitate the subsequent computation of answer-aware states. In Figure 2, (b) and (c) are identical, except (c) incorporates an additional vanilla BART encoder Lewis et al. (2020) to encode keywords. Although structurally similar, they are independent because of their distinct tasks. Specifically, (b) generates keywords based on the document and answer, while (c) leverages the keywords to pinpoint essential information snippets related to QA pairs in the document and generate multi-hop questions. Formalizing these processes:
@@ 4.1 Dataset
Table 1: Statistics of HotpotQA, where "Sam." denotes "Sample," "Qes." denotes "Question keyword," and "Doc." denotes "Document keyword."
@@ C.2 Results of Fine-tuned Large Language Models
It can be seen from Table 8 that in both the SF and Full settings, Llama’s performance on BLEU-4 and METEOR is inferior to that of the DPKG framework, while it performs better on ROUGE-L and BERTScore. Although Llama has a large number of parameters, it does not demonstrate stronger performance compared to the DPKG framework. On the one hand, this highlights the difficulty of the MQG task; on the other, it underscores the effectiveness of the dual-perspective keywords and the DPKG framework. These findings again suggest the challenges large models face in generating multi-hop questions — a problem that remains to be addressed in future research.
@@ Appendix E Generalization to Other Datasets
Table 12: Statistics of MusiQue-2hop, where "Sam." denotes "Sample," "Qes." denotes "Question keyword," and "Doc." denotes "Document keyword."
Model B4 MTR R-L